What are the atomic number element whose outermost element are represented by 3s1 and 2p3

Grade:11

1 Answers

jyoti bhatia
5 years ago
1s22s22p63s1: Atomic number of element is 11.
1s22s22p3: Atomic number of element is 7.

Completing the electron configuration of the element as 1s2 2s22p6 3s1.

∴ Number of electrons present in the atom of the element

= 2 + 2 + 6 + 1 = 11

∴ Atomic number of the element = 11(sodium).

Completing the electron configuration of the element as 1s2 2s22p3.

∴ Number of electrons present in the atom of the element = 2 + 2 + 3 = 7

∴ Atomic number of the element = 7(nitrogen).